OpenStudy (anonymous):

Which equation is correct? sec x° = opposite ÷ adjacent cot x° = opposite ÷ adjacent cosec x° = opposite ÷ adjacent sec x° = hypotenuse ÷ adjacent

OpenStudy (mathstudent55):

Did you ever hear of SOHCAHTOA?

OpenStudy (mathstudent55):

SOH: Sin = Opp/Hyp CAH: Cos = Adj/Hyp TOA: Tan = Opp/Adj Then you need to know this for sec, csc, and cot: sec = 1/cos csc = 1/sin cot = 1/tan